OpenStudy (anonymous):

The three lines y=2x,y=6-x,and y=-2 intersect to form a triangle. What are the coordinates of the vertices of the triangle?

OpenStudy (anonymous):

solve 3 sets y = 2x y = 6-x y = 2x y = -2 y = 6-x y = -2

OpenStudy (anonymous):

what do you mean by solve them?

OpenStudy (anonymous):

y = 2x y = 6-x can you find "y" and "x" that satisfy the above expressions?~

OpenStudy (anonymous):

how do i do that? i never learned this :S

OpenStudy (anonymous):

oh ok so you have y = 2x and y = 6-x, using y = 2x, you substitute it into y = 6-x y = 6-x substitute y = 2x 2x = 6-x + x both sides 3x = 6 divide by 3 both sides x = 2 with x = 2, substitute into y = 2x so you have y =4

OpenStudy (anonymous):

that's for the first set, one of the vertex of the triangle is (2,4) do similarly for the other 2 sets~

OpenStudy (anonymous):

thank you!!!